OpenStudy (anonymous):

Use the five steps for problem solving to answer the following question. Please show all of your work. The sum of the length and girth of a rectangular box must be smaller than 96 inches. If the girth is one half the length, what is the length? This is what I got. Is it right?? Length + girth = 96inches girth = 1/2 length (therefore, length = 2 girth) length + girth = 2 girth + 1 girth = 3 girth = 96 inches 3 girth = 96 inches => girth = 32 inches length = 2 girth = 64 inches

OpenStudy (stormfire1):

That looks like the right answer if you ignore the fact that it says it must be *smaller* than 96 inches :) Where x and y = length and girth respectively: x + y < 96 y = (1/2)x x + (1/2)x < 96 (3/2)x < 96 x < 64

OpenStudy (anonymous):

How about this let L = length; g=girth then L+g<96 now we are given that g=L/2, so substitute L+.5L<96 1.5L<96 L<64

OpenStudy (stormfire1):

Technically, I see more than one answer here. Since you know the length must be < 96, you could go with the length being 32 and the girth being 16...which answers the original question.

OpenStudy (stormfire1):

You could pretty must put any to values in there where one value is 2x the other and as long as the sum is < 96, it would be correct. My guess is that the teacher probably wants you to show the steps that led to you knowing that the x must be < 64

OpenStudy (anonymous):

yes he does.. and thanks so much for your help

OpenStudy (stormfire1):

That should have read: "You could pretty much put any two.." :) Np!